# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v1

@@ 4.2 Dash Phase
To leverage the identified LTSD, we directly learn the changes in their coordinates, denoted as $\Delta{\sigma}_{i}$ for the $i$-th $\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}$, to further fine-tune the model for downstream tasks. $\Delta{\sigma}_{i}$ is initialized as zero. Mathematically, this adjustment is represented by
@@
This equation forms the basis of LoRA-Dash, which aims to fully utilize the power of LTSD derived from LoRA. During training, $\mathbf{A}$, $\mathbf{B}$, and $\Delta{\sigma}_{i}$ are continuously updated. For more details on LoRA-Dash, please refer to Sec. D.
Figure 3: Framework of LoRA-Dash. LoRA-Dash comprises two phases: the pre-launch phase for identifying TSD, and the dash phase to maximize their potential.
@@ 5 How Much Can TSD Boost the Performance of LoRA-Dash?
To explore the performance gains of LoRA-Dash, both numerical and visual comparisons are employed. The numerical experiments focus on commonsense reasoning and natural language understanding tasks, assessing LoRA-Dash’s performance improvements in standard natural language processing metrics. Visual comparisons focuses on subject-driven generation task, involving using a text-to-image diffusion model to learn specific concepts (i.e., task-specific concept), providing a clear, qualitative and intuitive understanding of the impact of TSD on performance. The implementation details are shown in Sec. E.1. The experiment details are shown in Secs. E.4-E.6.
@@ 5.1 Numerical Results
Robustness to Parameter Budget: (When fine-tuning LLaMA-7B) LoRA shows enhanced performance with an increased parameter budget, indicating its dependence on larger rank sizes. This highlights its effectiveness in harnessing the full potential of TSD, even in constrained settings. Conversely, LoRA-Dash maintains strong performance even under limited parameter conditions, which suggests that TSD are crucial for maximizing fine-tuning efficiency.
@@ 5.2 Visual Results
Moreover, the effectiveness of TSD is particularly evident in the response of LoRA-Dash to given prompts. TSD effectively encapsulate the specific information of each prompt, enabling LoRA-Dash to precisely grasp the semantic essence of prompts such as sea beach or wheat field. This capability allows LoRA-Dash to adeptly render complex themes and nuanced details within the images.
@@ 6 Understanding LoRA-Dash
In this section, we conduct extensive experiments to better understand LoRA-Dash. We first explore the relationships between the LTSD, the final directions after the dash phase, and TSD to determine whether LoRA-Dash can amplify the information of TSD (Sec. 6.1) and, if so, to what extent (Sec. 6.2). We then compare the effectiveness of different directions, such as the top singular directions (Sec. 6.3), and investigate the impact of the length of the pre-launch phase and the number of directions in the dash phase on performance (Sec. 6.4). Finally, we verify whether the information in TSD can help enhance the effectiveness of other methods (Sec. 6.5). We have also provided extensive discussions in the supplementary material, and we strongly encourage readers to consult it for a more comprehensive understanding of LoRA-Dash.
@@ 6.1 How Does LoRA-Dash Determine and Align with Task-specific Directions
We examine the alignment between the LTSD identified during the pre-launch phase, the directions with highest change rates pinpointed by the final trained $\Delta\mathbf{W}_{all}$ of LoRA-Dash (denoted as “DTSD”, Delta TSD), and the true TSD derived from $\mathbf{W}^{*}$. Following the settings in Secs. 3.4 and E.3, we report the percentage that:
  • How many of the top 4 DTSD directions are contained in $s$ LTSD [^4] Here since $s=8$, the setting is equal to that in Sec. E.3.. We denote this percentage as “DTSD $\cap$ LTSD ” for further convenience.
  • How many of the top 4 TSD are contained in $s$ LTSD. Denote this as “TSD $\cap$ LTSD”
  • How many of the top 4 TSD are in $s$ DTSD directions. Denote this as “TSD $\cap$ DTSD”.
The first is to explore the consistency between the directions selected during the pre-launch phase of LoRA-Dash and those showing the greatest change after the final training, and the second and third aim to validate the consistency between the directions chosen during the LoRA-Dash pre-launch phase and the final trained with the true TSD. We conducted tests on LLaMA-7B Touvron et al. (2023a) and DeBERTaV3-large He et al. (2021b), each with three ranks of LoRA-Dash, and the results are shown in Tables. 3 and 20-21.
Table 3: Alignment on the directions of LoRA-Dash and TSD. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules. More results are shown in Tables. 20-21.
@@
DTSD $\cap$ LTSD 0.75 1.00 0.75 0.77 1.00 1.00 1.00 0.98
TSD $\cap$ LTSD 0.75 0.75 1.00 0.74 1.00 1.00 1.00 0.84
TSD $\cap$ DTSD 1.00 0.75 1.00 0.82 1.00 1.00 1.00 0.89
It is obvious that LoRA-Dash indeed captures a great proportion of TSD’s information. Moreover, we can also obtain two significant findings: First, the LTSD and DTSD of LoRA-Dash align closely, which indicates that the directions amplified by LoRA-Dash finally are essentially consistent with those selected during the pre-launch phase. Therefore, since LTSD contains substantial information about TSD, it is not surprising that the final DTSD consistently include information about the true TSD. Second, when compared with the results in Fig. 7 where the end of each line indicates the final percentage of TSD $\cap$ DTSD of LoRA, it becomes evident that LoRA-Dash’s trained weights capture a greater proportion of TSD information than vanilla LoRA.
@@ 6.2 How much does LoRA-Dash Amplify Task-specific Directions?
Since LoRA-Dash captures the information of TSD, we further explore to what extent LoRA-Dash can amplify the information of these directions. To quantify this amplification, we define an amplification factor that measures how significantly LoRA-Dash enhances the task-specific features. The details are shown in Sec. E.7.
@@
We conduct this experiment on LLaMA-7B with three ranks. We also test the separate impacts of $\Delta\mathbf{W}_{AB}$ and $\Delta\mathbf{W}_{dash}$ on the amplification. The results shown in Table. 4 demonstrate that LoRA-Dash significantly enhances the features associated with LTSD under all test conditions. Notably, the feature amplification contributed by $\Delta\mathbf{W}_{dash}$ accounts for the majority of the overall enhancement, whereas the amplification from $\Delta\mathbf{W}_{AB}$ is relatively minor in comparison. LoRA-Dash effectively learns and amplifies the majority of TSD information with only $s$ parameters (i.e., $\Delta\sigma_{i}$ in $\Delta\mathbf{W}_{dash}$), reducing the learning burden on $\Delta\mathbf{W}_{AB}$ during fine-tuning. This efficiency also allows $\Delta\mathbf{W}_{AB}$ to potentially focus on learning other significant features, further optimizing the model’s performance on downstream tasks. We strongly encourage readers to refer to Sec. E.7 for a deeper discussion on this topic.
@@ 6.3 The Effectiveness of TSD Compared with Other Core Directions
We investigated the effects of TSD during the pre-launch phase compared to other locations, such as selecting top, bottom or random core directions. The results when fine-tuning LLAMA3-8B, as shown in Fig. 5(a), clearly indicate that selecting TSD yields the best performance, followed by dash-bottom, with top selections performing the worst. This is not difficult to understand: directly dashing TSD maximizes the ability of $\mathbf{W}$ to adapt to downstream tasks. Since TSD are generally located towards the end of the spectrum, choosing directions from the bottom naturally results in better outcomes than selecting from the top. We have a more thorough discussion in Sec. F.3.
@@ 6.4 The influence of the Hyper-parameters in LoRA-Dash
We evaluate the influence of two hyper-parameters $t$ and $s$ in LoRA-Dash when fine-tuning LLAMA-7B, and the results are illustrated in Fig. 5(b)-(c). It can be observed that starting the dash phase after a longer pre-launch period tends to yield slightly inferior results compared to starting it earlier. This could be attributed to the fact that entering the dash phase earlier allows for more extensive utilization of TSD, potentially enhancing the adaptation to downstream tasks more effectively. Additionally, a larger value of $s$, representing the number of directions used in the dash phase, also impacts performance. Based on our analysis in Sec. E.7, we suspect that the inclusion of some irrelevant directions might introduce noise into the training process (i.e., destabilizing model convergence), leading to diminished performance.
@@ 6.5 Can TSD Enhance the Performance of Other Methods?
To investigate whether TSD are beneficial for other PEFT methods, we extended our experiments to include AdaLoRA Zhang et al. (2022) and FLoRA Si et al. (2024a), assessing the impact of leveraging TSD within these frameworks. The results, as shown in Table. 5, clearly indicate that incorporating TSD significantly enhances the performance of these methods as well. This underscores the pivotal role of TSD in optimizing model behavior for downstream tasks, highlighting their universal applicability across various PEFT strategies.
@@ 7 Conclusion
In this paper, we revisit LoRA’s exploration of task-specific directions (TSD), highlighting its inadequacies in understanding TSD. We then provide a precise definition of TSD and delve into their properties to better understand their role in fine-tuning large language models. Building on this foundational knowledge, we introduce LoRA-Dash, designed to fully unleash the potential of TSD. Through comprehensive experiments and in-depth analysis, we demonstrate the significant advantages of LoRA-Dash over conventional methods. Our findings not only validate the effectiveness of LoRA-Dash but also highlight the crucial importance of TSD in achieving superior task-specific performance. By pushing the boundaries of parameter-efficient fine-tuning, we aim to inspire continued research and development in this vibrant area, transforming practices across diverse applications in natural language processing and beyond.
Appendix A Background
A.1 Parameter Efficient Fine-tuning
A.2 Recent Research on Task-specific Directions
LoRA highlights that $\Delta\mathbf{W}$ encompasses directions within $\mathbf{W}$ that are deemed insignificant for initial tasks but are crucial for specific downstream tasks, which LoRA refers to as “task-specific” directions. During our survey of existing works, we discovered that several methods, although not explicitly stated, appear to adjust these directions. A common approach involves decomposing $\mathbf{W}$ using singular vector decomposition (SVD) and then adjusting either the top singular directions Meng et al. (2024); Zhang & Pilanci (2024), the bottom singular directions Wang et al. (2024), or selecting singular directions randomly Hameed et al. (2024). However, we argue that this kind of practice is unreasonable and can not truly identify the task-specific directions. According to the analysis in Sec. 6.1, task-specific directions are not simply identified by being at the top, bottom, or through random selection. Moreover, even if a direction selected by such a simple method is genuinely task-specific, altering it might cause it to lose its task-specific quality.
Appendix B Structure and Contributions of This Paper
@@
  • We begin by reviewing the description and exploration of TSD within the LoRA framework, examining its properties. Our analysis reveals a clear contradiction in LoRA’s conclusions regarding TSD, stemming from an ambiguous definition of TSD within the framework (Sec. 2).
  • To address this, we propose a novel framework that clearly defines what TSD are (Sec. 3.2) and further explores their properties (Sec. 3.3). Additionally, we highlight the challenges of applying TSD in fine-tuning processes (Sec. 3.4).
  • To further unleash the potential of TSD, we propose a new method, LoRA-Dash, which leverages TSD to enhance performance on downstream tasks (Sec. 4). We first introduce how LoRA-Dash identifies TSD (Sec. 4.1), followed by an explanation of how it leverages this information (Sec. 4.2).
@@
  • We propose a framework, which clearly defines TSD, and offers foundations for further analysis.
  • We propose a novel PEFT method LoRA-Dash, which identified TSD and leverage the potential of them to enhance performance on downstream tasks.
Appendix C Details on Task-specific Directions
Figure 6: Pseudo codes of LoRA-Dash. LoRA is capable of identifying TSD, but it does not utilize the corresponding information. LoRA-Dash builds on this by actively leveraging the power of these identified TSD to enhance model performance significantly.
C.1 How Do Task-specific Directions Arise in LoRA?
@@
Additionally, LoRA introduces the “feature amplification factor” to measure the extent of feature enhancement, defined as ${\|\Delta\mathbf{W}\|_{F}}/{\|\mathbf{U}^{\mathsf{T}}\mathbf{W}\mathbf{V}\|_{F}}$. The factor is significantly higher when $\mathbf{U}$ and $\mathbf{V}$ are top singular vectors derived from $\Delta\mathbf{W}$ compared to when they are those derived from $\mathbf{W}$, suggesting that $\Delta\mathbf{W}$ only boosts directions that are not emphasized in $\mathbf{W}$ (conclusion 2). Moreover, a larger $r$ ($r=64$) yields a much lower amplification factor than a smaller $r$ ($r=4$), implying that the number of “task-specific directions” is small (conclusion 1). Through extensive experiments across various ranks and layers, LoRA determines that $\Delta\mathbf{W}$ potentially amplifies important directions for specific downstream tasks that were learned but not emphasized during general pre-training (conclusion 3), and refers to them as “task-specific” directions (TSD).
@@
C.2 Why should TSD be Established on $\mathbf{W}$?

### v2

@@ 4.2 Dash Phase
To leverage the identified LTSDs, we directly learn the changes in their coordinates, denoted as $\Delta{\sigma}_{i}$ for the $i$-th $\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}$, to further fine-tune the model for downstream tasks. $\Delta{\sigma}_{i}$ is initialized as zero. Mathematically, this adjustment is represented by
@@
This equation forms the basis of LoRA-Dash, which aims to fully utilize the power of LTSDs derived from LoRA. During training, $\mathbf{A}$, $\mathbf{B}$, and $\Delta{\sigma}_{i}$ are continuously updated. For more details on LoRA-Dash, please refer to Sec. C.
Figure 3: Framework of LoRA-Dash. LoRA-Dash comprises two phases: the pre-launch phase for identifying TSDs, and the dash phase to maximize their potential.
@@ 5 How Much Can TSD Boost the Performance of LoRA-Dash?
To explore the performance gains of LoRA-Dash, both numerical and visual comparisons are employed. The numerical experiments focus on commonsense reasoning and natural language understanding tasks, assessing LoRA-Dash’s performance improvements in standard natural language processing metrics. Visual comparisons focuses on subject-driven generation task, involving using a text-to-image diffusion model to learn specific concepts (i.e., task-specific concept), providing a clear, qualitative and intuitive understanding of the impact of TSDs on performance. The implementation details are shown in Sec. D.1. The experiment details are shown in Secs. D.4-D.6.
@@ 5.1 Numerical Results
Robustness to Parameter Budget: (When fine-tuning LLaMA-7B) LoRA shows enhanced performance with an increased parameter budget, indicating its dependence on larger rank sizes. This highlights its effectiveness in harnessing the full potential of TSDs, even in constrained settings. Conversely, LoRA-Dash maintains strong performance even under limited parameter conditions, which suggests that TSDs are crucial for maximizing fine-tuning efficiency.
@@ 5.2 Visual Results
Moreover, the effectiveness of TSDs is particularly evident in the response of LoRA-Dash to given prompts. TSDs effectively encapsulate the specific information of each prompt, enabling LoRA-Dash to precisely grasp the semantic essence of prompts such as sea beach or wheat field. This capability allows LoRA-Dash to adeptly render complex themes and nuanced details within the images.
@@ 6 Understanding LoRA-Dash
In this section, we conduct extensive experiments to better understand LoRA-Dash. We first explore the relationships between the LTSDs, the final directions after the dash phase, and TSDs to determine whether LoRA-Dash can amplify the information of TSDs (Sec. 6.1) and, if so, to what extent (Sec. 6.2). We then compare the effectiveness of different directions, such as the top singular directions (Sec. 6.3), and investigate the impact of the length of the pre-launch phase and the number of directions in the dash phase on performance (Sec. 6.4). Finally, we verify whether the information in TSDs can help enhance the effectiveness of other methods (Sec. 6.5). We have also provided extensive discussions in the supplementary material, and we strongly encourage readers to consult it for a more comprehensive understanding of LoRA-Dash.
@@ 6.1 How Does LoRA-Dash Determine and Align with Task-specific Directions
We examine the alignment between the LTSDs identified during the pre-launch phase, the directions with highest change rates pinpointed by the final trained $\Delta\mathbf{W}_{all}$ of LoRA-Dash (denoted as “DTSDs”, Delta TSDs), and the true TSDs derived from $\mathbf{W}^{*}$. Following the settings in Secs. 3.4 and D.3, we report the percentage that:
  • How many of the top 4 DTSDs directions are contained in $s$ LTSDs [^3] Here since $s=8$, the setting is equal to that in Sec. D.3.. We denote this percentage as “DTSDs $\cap$ LTSDs ” for further convenience.
  • How many of the top 4 TSDs are contained in $s$ LTSDs. Denote this as “TSDs $\cap$ LTSDs”
  • How many of the top 4 TSDs are in $s$ DTSDs. Denote this as “TSDs $\cap$ DTSDs”.
The first is to explore the consistency between the directions selected during the pre-launch phase of LoRA-Dash and those showing the greatest change after the final training, and the second and third aim to validate the consistency between the directions chosen during the LoRA-Dash pre-launch phase and the final trained with the true TSDs. We conducted tests on LLaMA-7B Touvron et al. (2023a) and DeBERTaV3-large He et al. (2021b), each with three ranks of LoRA-Dash, and the results are shown in Tables. 3 and 20-21.
Table 3: Alignment on the directions of LoRA-Dash and TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules. More results are shown in Tables. 20-21.
@@
DTSDs $\cap$ LTSDs 0.75 1.00 0.75 0.77 1.00 1.00 1.00 0.98
TSDs $\cap$ LTSDs 0.75 0.75 1.00 0.74 1.00 1.00 1.00 0.84
TSDs $\cap$ DTSDs 1.00 0.75 1.00 0.82 1.00 1.00 1.00 0.89
It is obvious that LoRA-Dash indeed captures a great proportion of TSDs’ information. Moreover, we can also obtain two significant findings: First, the LTSDs and DTSDs of LoRA-Dash align closely, which indicates that the directions amplified by LoRA-Dash finally are essentially consistent with those selected during the pre-launch phase. Therefore, since LTSDs contain substantial information about TSDs, it is not surprising that the final DTSDs consistently include information about the true TSDs. Second, when compared with the results in Fig. 7 where the end of each line indicates the final percentage of TSDs $\cap$ DTSDs of LoRA, it becomes evident that LoRA-Dash’s trained weights capture a greater proportion of TSD information than vanilla LoRA.
@@ 6.2 How much does LoRA-Dash Amplify Task-specific Directions?
Since LoRA-Dash captures the information of TSDs, we further explore to what extent LoRA-Dash can amplify the information of these directions. To quantify this amplification, we define an amplification factor that measures how significantly LoRA-Dash enhances the task-specific features. The details are shown in Sec. D.7.
@@
We conduct this experiment on LLaMA-7B with three ranks. We also test the separate impacts of $\Delta\mathbf{W}_{AB}$ and $\Delta\mathbf{W}_{dash}$ on the amplification. The results shown in Table. 4 demonstrate that LoRA-Dash significantly enhances the features associated with LTSDs under all test conditions. Notably, the feature amplification contributed by $\Delta\mathbf{W}_{dash}$ accounts for the majority of the overall enhancement, whereas the amplification from $\Delta\mathbf{W}_{AB}$ is relatively minor in comparison. LoRA-Dash effectively learns and amplifies the majority of TSD information with only $s$ parameters (i.e., $\Delta\sigma_{i}$ in $\Delta\mathbf{W}_{dash}$), reducing the learning burden on $\Delta\mathbf{W}_{AB}$ during fine-tuning. This efficiency also allows $\Delta\mathbf{W}_{AB}$ to potentially focus on learning other significant features, further optimizing the model’s performance on downstream tasks. We strongly encourage readers to refer to Sec. D.7 for a deeper discussion on this topic.
@@ 6.3 The Effectiveness of TSD Compared with Other Core Directions
We investigated the effects of TSDs during the pre-launch phase compared to other locations, such as selecting top, bottom or random core directions. The results when fine-tuning LLAMA3-8B, as shown in Fig. 5(a), clearly indicate that selecting TSDs yields the best performance, followed by dash-bottom, with top selections performing the worst. This is not difficult to understand: directly dashing TSDs maximizes the ability of $\mathbf{W}$ to adapt to downstream tasks. Since TSDs are generally located towards the end of the spectrum, choosing directions from the bottom naturally results in better outcomes than selecting from the top. We have a more thorough discussion in Sec. E.3.
@@ 6.4 The influence of the Hyper-parameters in LoRA-Dash
We evaluate the influence of two hyper-parameters $t$ and $s$ in LoRA-Dash when fine-tuning LLAMA-7B, and the results are illustrated in Fig. 5(b)-(c). It can be observed that starting the dash phase after a longer pre-launch period tends to yield slightly inferior results compared to starting it earlier. This could be attributed to the fact that entering the dash phase earlier allows for more extensive utilization of TSDs, potentially enhancing the adaptation to downstream tasks more effectively. Additionally, a larger value of $s$, representing the number of directions used in the dash phase, also impacts performance. Based on our analysis in Sec. D.7, we suspect that the inclusion of some irrelevant directions might introduce noise into the training process (i.e., destabilizing model convergence), leading to diminished performance.
@@ 6.5 Can TSD Enhance the Performance of Other Methods?
To investigate whether TSDs are beneficial for other PEFT methods, we extended our experiments to include AdaLoRA Zhang et al. (2022) and FLoRA Si et al. (2024a), assessing the impact of leveraging TSDs within these frameworks. The results, as shown in Table. 5, clearly indicate that incorporating TSDs significantly enhances the performance of these methods as well. This underscores the pivotal role of TSDs in optimizing model behavior for downstream tasks, highlighting their universal applicability across various PEFT strategies.
@@ 7 Conclusion
In this paper, we revisit LoRA’s exploration of task-specific directions (TSDs), highlighting its inadequacies in understanding TSDs. We then provide a precise definition of TSD and delve into their properties to better understand their role in fine-tuning large language models. Building on this foundational knowledge, we introduce LoRA-Dash, designed to fully unleash the potential of TSDs. Through comprehensive experiments and in-depth analysis, we demonstrate the significant advantages of LoRA-Dash over conventional methods. Our findings not only validate the effectiveness of LoRA-Dash but also highlight the crucial importance of TSDs in achieving superior task-specific performance. By pushing the boundaries of parameter-efficient fine-tuning, we aim to inspire continued research and development in this vibrant area, transforming practices across diverse applications in natural language processing and beyond.
Appendix A Structure and Contributions of This Paper
@@
  • We begin by reviewing the description and exploration of TSDs within the LoRA framework. Our analysis reveals the significance of the clear definition of TSDs within PEFT (Sec. 2).
  • We propose a novel framework that clearly defines what TSDs are (Sec. 3.2) and further explores their properties (Sec. 3.3). Additionally, we highlight the challenges of applying TSDs in fine-tuning processes (Sec. 3.4).
  • To further unleash the potential of TSDs, we propose a new method, LoRA-Dash, which leverages TSDs to enhance performance on downstream tasks (Sec. 4). We first introduce how LoRA-Dash identifies TSDs (Sec. 4.1), followed by an explanation of how it leverages this information (Sec. 4.2).
@@
  • We propose a framework, which clearly defines TSDs, and offers foundations for further analysis.
  • We propose a novel PEFT method LoRA-Dash, which identified TSDs and leverage the potential of them to enhance performance on downstream tasks.
Appendix B Details on Task-specific Directions
Figure 6: Pseudo codes of LoRA-Dash. LoRA is capable of identifying TSDs, but it does not utilize the corresponding information. LoRA-Dash builds on this by actively leveraging the power of these identified TSDs to enhance model performance significantly.
B.1 How Do Task-specific Directions Arise in LoRA?
@@
Additionally, LoRA introduces the “feature amplification factor” to measure the extent of feature enhancement, defined as ${\|\Delta\mathbf{W}\|_{F}}/{\|\mathbf{U}^{\mathsf{T}}\mathbf{W}\mathbf{V}\|_{F}}$. The factor is significantly higher when $\mathbf{U}$ and $\mathbf{V}$ are top singular vectors derived from $\Delta\mathbf{W}$ compared to when they are those derived from $\mathbf{W}$, suggesting that $\Delta\mathbf{W}$ only boosts directions that are not emphasized in $\mathbf{W}$. Moreover, a larger $r$ ($r=64$) yields a much lower amplification factor than a smaller $r$ ($r=4$), implying that the number of “task-specific directions” is small. Through extensive experiments across various ranks and layers, LoRA determines that $\Delta\mathbf{W}$ potentially amplifies important directions for specific downstream tasks that were learned but not emphasized during general pre-training, and refers to them as “task-specific” directions (TSDs).
LoRA has drawn three conclusions related to TSDs:
  1. 1.
    TSDs are the directions of top singular vectors derived from the learned $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$. TSDs consistently exhibit significant overlap across different $r$ settings used in the learning of $\Delta\mathbf{W}$, where $r$ is the rank configuration of $\Delta\mathbf{W}$.
  2. 2.
    TSDs are not the directions of top singular vectors derived from the pretrained weights $\mathbf{W}$.
  3. 3.
    TSDs are certain directions that have already been learned by $\mathbf{W}$ but were not emphasized.
However, only the first conclusion of LoRA already presents significant contradictions. For a specific task, TSDs should indeed remain consistent and not depend on learned $\Delta\mathbf{W}$. Moreover, revisiting the first and the third conclusions, we could deduce that the top singular directions of $\Delta\mathbf{W}$ are among the directions of $\mathbf{W}$. Nonetheless, in practice, the singular directions of $\Delta\mathbf{W}$ are unlikely to align exactly with any specific directions in $\mathbf{W}$. As shown in our paper, although LoRA highlights the importance of TSDs for downstream tasks, it struggles with a clear definition of TSDs which leads to contradictions, and does not effectively outline how they can be leveraged, leading to potential under-utilization in practical applications
@@
B.2 Why should TSD be Established on $\mathbf{W}$?